If you have 2 vector quantities, one that is 4m, and one that is 6m, use vector math to determine the length of the resultant vector
1
Expert's answer
2019-09-17T12:08:56-0400

We cannot find vector sum of these two vectors because we didn't know the angle between them.

The length of the resultant vector:


"R=\sqrt{4^2+6^2+2(4)(6)\cos{\alpha}}"

"R=\sqrt{52+48\cos{\alpha}}"